# Stabilizing Reinforcement Learning for Honesty Alignment in Language Models on Deductive Reasoning

### v1

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has recently emerged as a promising framework for aligning language models with complex reasoning objectives. However, most existing methods optimize only for final task outcomes, leaving models vulnerable to collapse when negative rewards dominate early training. This challenge is especially pronounced in honesty alignment, where models must not only solve answerable queries but also identify when conclusions cannot be drawn from the given premises. Deductive reasoning provides an ideal testbed because it isolates reasoning capability from reliance on external factual knowledge. To investigate honesty alignment, we curate two multi-step deductive reasoning datasets from graph structures, one for linear algebra and one for logical inference, and introduce unanswerable cases by randomly perturbing an edge in half of the instances. We find that GRPO, with or without supervised fine tuning initialization, struggles on these tasks. Through extensive experiments across three models, we evaluate stabilization strategies and show that while curriculum learning provides some benefit, it requires careful design in distribution datasets with controllable difficulty. To address these limitations, we propose Anchor, a reinforcement learning method that injects ground truth trajectories into rollouts, preventing early training collapse. Our results demonstrate that this method stabilizes learning and significantly improves the overall reasoning performance, underscoring the importance of training dynamics for enabling reliable deductive reasoning in aligned language models.
@@ 1 Introduction
While large language models (LLMs) have demonstrated remarkable capabilities in knowledge-intensive tasks and complex reasoning, their increasing deployment in real-world applications introduces critical safety considerations (betley2025emergent; raza2025industrial; bengio2025superintelligent; cloud2025subliminal). For these models to be deployed reliably, it is not sufficient for them to be merely helpful and harmless: they must also be honest (askell2021general; greenblatt2024alignmentfakinglargelanguage; sheshadri2025languagemodelsfakealignment). This means a model should not just avoid fabricating information, but should actively acknowledge the boundaries of its competence rather than misleading users with unwarranted confidence (jiang2021can; yin2023large; mohri2024language; kalai2025languagemodelshallucinate). Unfortunately, current training paradigms and evaluation benchmarks often fail to incentivize or measure this crucial behavior (li2025elicitinglanguagemodelbehaviors; chowdhury2025surfacing).
@@
Supervised fine-tuning (SFT) (wu2025generalization) has proven highly effective at quickly aligning models to desired behaviors. However, it is well noted that SFT tends to overfit to demonstrations and struggles to generalize beyond the dataset distribution, especially for reasoning tasks. Reinforcement learning methods such as Group Relative Policy Optimization (GRPO) (shao2024deepseekmath), which compares multiple rollouts of the same query to assign relative advantages, instead optimize for final verifiable outcomes. Yet, prior analyses have shown that when all rollouts are incorrect and receive identical rewards, their relative advantages collapse to zero, leading to several issues including vanishing gradients and reinforcing dishonest overconfidence (liu2025understanding; yu2025dapo; zheng2025group). For honesty alignment, the challenge lies not only in training but also in evaluation. Existing benchmarks often conflate two distinct sources of failure: factual inaccuracy versus failure to acknowledge unanswerability, making it difficult to isolate and measure honesty in reasoning (joshi2017triviaqa; kwiatkowski-etal-2019-natural; li2023halueval; niu2023ragtruth; guan2024hallusionbench). This motivates the need for new datasets and methods that specifically target honest reasoning. [^1] For a comprehensive discussion of related work, see Appendix A.
@@
To study this problem in a controlled setting, we focus on honesty alignment in deductive reasoning, a reasoning paradigm where the answerability of a conclusion depends solely on whether it can be derived from the given premises in the prompt. This property eliminates dependence on external knowledge, cleanly separating reasoning ability from factual recall. To this end, we construct two multi-step reasoning datasets in which each query is either answerable or unanswerable given the premises, providing a precise testbed for evaluating whether models can reason honestly and recognize when valid reasoning paths exist and when they do not. The first dataset, GraphLA, is grounded in linear algebra: queries correspond to solving systems of equations along reasoning paths, while unanswerable cases are created by perturbing the system so that no valid solution path exists. The second dataset, GraphLI, is based on logical inference: queries test whether a conclusion follows from composed chains of implications, with unanswerable instances generated by removing or altering key premises or conclusions. Based on these datasets, we investigate the following two central research questions:
@@
For evaluation (RQ1), we show that across three model scales, performance on our benchmarks sharply declines as reasoning depth and problem size increase. Models struggle not only to follow reasoning chains but also to refrain from producing unwarranted conclusions when no valid derivation exists, revealing a lack of honest reasoning. For training (RQ2), we find that standard SFT and GRPO fail to overcome these challenges. Curriculum learning, when easy datasets are carefully and properly constructed, achieves strong performance but remains fragile and highly sensitive to difficulty calibration. In contrast, Anchor consistently stabilizes reinforcement learning, achieving robust performance on both answerable and unanswerable queries. When paired with curriculum learning, Anchor provides further gains, underscoring its strength in guiding models toward stable and honest reasoning.
@@
    We formalize honesty alignment as the ability to abstain on unanswerable queries, and introduce two deductive reasoning datasets, GraphLA and GraphLI, that balance answerable and unanswerable cases.
@@ 3 RQ1 (Motivation): How Do Untrained Models Perform on Reasoning Tasks of Varying Deductive Difficulty?
Experiment Setup
We evaluate three open-sourced models: Qwen-2.5-3B-Instruct, Qwen-3-0.6B, and Qwen-3-1.7B. Additional details are provided in § D.1.
@@ 4 RQ2 (Core Problem): How Can Training Equip Models with Honest Reasoning Abilities?
4.1 Preliminaries
Group Relative Policy Optimization (GRPO)
GRPO (shao2024deepseekmath) computes the relative advantage of each response within a group of responses to the same query. The corresponding policy gradient is
Supervised Fine-tuning (SFT)
The objective of SFT is to maximize the likelihood of ground-truth responses sampled from a supervised dataset. The SFT objective gradient is simply the logarithmic likelihood gradient on the supervised dataset, with no advantage weighting:
@@ 5 Experiments
We experiment with the following baselines and approaches: Chain-of-Thought (CoT) prompting (wei2022chain), supervised fine-tuning (SFT), GRPO (shao2024deepseekmath), SFT+GRPO (where SFT is used as a cold start followed by GRPO), and Easy-to-Hard curriculum learning (where GRPO is first trained on an easier dataset and then on the target dataset).
@@
For GraphLI, we fix the reasoning depth to $k=15$ and the number of irrelevant edges to $|E_{\text{irr}}|=5$. For each configuration, 3 examples are sampled for instantiation. The dataset is split into 5316 training (2702 answerable, 2614 unanswerable), 300 validation (143 answerable, 157 unanswerable), and 300 test (155 answerable, 145 unanswerable) examples. For the easy dataset in curriculum learning, we reduce the reasoning depth to $k\in\{2,3,4,5\}$.
We evaluate on Qwen-2.5-3B-Instruct, Qwen-3-0.6B, and Qwen-3-1.7B. As evaluation metrics, we report overall accuracy, accuracy on the unanswerable subset, and accuracy on the answerable subset. Additional details on the experiment setup are provided in Appendix G.
@@ 7 Conclusion
We investigated the challenge of aligning reasoning language models with honesty, focusing on tasks that require both solving answerable queries and abstaining on unanswerable ones. Our analysis showed that existing approaches such as SFT and GRPO either fail to provide reliable learning signals or collapse when faced with uniformly negative rewards. We proposed Anchor, a ground-truth–injected reinforcement learning method that stabilizes training by ensuring positive reference signals during rollouts. Across both the GraphLA and GraphLI datasets and multiple models, Anchor consistently outperformed baselines and proved robust where curriculum learning was fragile. Moreover, Anchor integrates seamlessly with Easy-to-Hard training, yielding further gains. These results highlight the importance of stabilizing reinforcement learning dynamics for enabling honest and reliable reasoning in language models.
@@
We have taken several steps to ensure the reproducibility of our work. A detailed description of dataset construction and statistics is provided in §§ 3 and 5, along with complete methodology details in § 4. Parameters and hyperparameters used for training and evaluation are documented in §§ 5, D.1 and G. Theoretical guarantees for our proposed method are supported by a formal proof of Proposition 1 in Appendix F. To further facilitate reproducibility, we will release both the datasets and source code upon acceptance.
Appendix A Related Work
A.1 Stabilizing Reinforcement Learning
SFT (i.e., behavior cloning) can be viewed as a special case of RL (wu2025generalization), where the policy is trained to imitate expert trajectories without any exploration. Early RL methods sought to overcome this limitation by explicitly introducing exploration. REINFORCE (williams1992simple) provided the first unbiased policy gradient estimator from sampled rollouts, but suffered from high variance. Generalized Advantage Estimation (GAE) (schulman2015high) mitigated this issue by balancing bias and variance in advantage estimation, thereby enabling more stable credit assignment. Building on these ideas, Proximal Policy Optimization (PPO) (schulman2017proximal) introduced a clipped surrogate objective to further stabilize training. More recently, GRPO (shao2024deepseekmath) also employs a clipped objective, but removes the value function and estimates advantages in a group-relative manner. GRPO is the first approach to compute the mean loss across tokens within each sequence before averaging across samples. Despite these innovations, GRPO has been shown to suffer from several limitations, including biased optimization toward longer responses and against instances that are either too easy or too hard (liu2025understanding), entropy collapse under poor exploration (yu2025dapo), and noisy training gradients due to token-level importance ratios (zheng2025group).
Our method, Anchor, specifically addresses the limitation that GRPO is poorly optimized for overly difficult queries, by introducing an SFT-like term into its gradient update. Several related approaches share similar motivations. Compared to it and SFT, Anchor avoids overfitting to demonstrations by combining them with policy-generated rollouts and providing a safeguard against vanishing gradients in GRPO when all rollouts fail. Deep Q-learning from Demonstrations (hester2018deep) incorporates demonstrations by storing them in a replay buffer and sampling them stochastically, whereas Anchor deterministically injects demonstrations into every rollout group at training time. This makes Anchor particularly well-suited for reasoning tasks where complete ground-truth trajectories are available but exploration often collapses without anchoring. Proximal Supervised Fine-Tuning (PSFT) (zhu2025proximal) tackles the overfitting issue of SFT by constraining policy updates during imitation and weighting equally. In contrast, Anchor produces a relative training signal by anchoring exploration with demonstrations in the same training loop, thereby aligning supervised and reinforcement signals more dynamically. As a concurrent line of work, chen2025beyond unify SFT and RL via bilevel optimization, where supervised fine-tuning serves as the upper-level objective that guides reinforcement learning at the lower level to maximize their “cooperative gain.” The strength of Anchor lies in its robustness and guaranteed learning progress, even when exploration fails (as is often the case in our experiments), whereas bilevel optimization may still stall when RL rollouts provide no meaningful reward signal. Several other recent studies have explored integrating SFT with RL. However, most either lack access to a verifiable ground-truth trajectory for guidance (yan2025learning; liu2025ghpo; huang2025blending; nath2025adaptive), or do not directly address the stability of policy gradient updates, leaving them still susceptible to zero-variance issues at certain training stages (wu2025thought; chen2025step; zhang2025policy).
@@ A.1 Stabilizing Reinforcement Learning
A.2 Honesty Alignment
askell2021general characterize AI alignment in terms of the “HHH” principles: helpful, honest, and harmless. Honesty requires that a model faithfully express its uncertainty, avoid misleading users, and acknowledge the limits of its own knowledge and capabilities. A broader notion of honesty also encompasses conflicts between generated outputs and ground truth, including factual correctness (allen2023physics; allen2023physics2; yang2024alignment) and proper uncertainty estimation or calibration (lin2022teaching; xiong2023can; geng-etal-2024-survey). Hallucination is an even broader phenomenon, where LLMs generate content that deviates from the input, contradicts previously generated context, or misaligns with established world knowledge (zhang2025siren). Honesty can be seen as a specific dimension of hallucination: the ability of a model to explicitly acknowledge when it is incapable of solving a problem, rather than producing an unwarranted conclusion.
Several benchmarks have been proposed to evaluate honesty (joshi2017triviaqa; kwiatkowski-etal-2019-natural; li2023halueval; niu2023ragtruth; guan2024hallusionbench). However, most knowledge-based benchmarks conflate two distinct sources of failure: (i) the model’s stored knowledge differs from the ground truth, which can in principle be mitigated by providing correct knowledge during training (wang2023survey; augenstein2024factuality); and (ii) the model fails to acknowledge that, given its own knowledge, it cannot solve the problem (yin2023large; ouyang2025treecut). This conflation makes it difficult to isolate and study honesty alignment in the strict sense of (ii). To our knowledge, kirichenko2025abstentionbench is the only recent work that directly addresses this dimension of honesty. Our datasets complement this line of research by targeting deductive reasoning without requiring external knowledge, a setting that naturally separates (ii) from (i). Finally, kalai2025language argue that hallucinations are a systemic outcome of current training and evaluation setups, and propose reforming evaluation standards so that admitting “I don’t know” is treated as a strength rather than a weakness. Our work aligns with this perspective by explicitly focusing on honest reasoning in tasks where models must distinguish between answerable and unanswerable queries.
@@ E.1 GRPO
All tokens within a rollout $y_{i}$ share the same normalized advantage $\hat{A}_{i}$.
@@ Appendix G Experiment Setup
Appendix H Clarifications on LLM Usage

### v2

# Toward Honest Language Models for Deductive Reasoning

@@ Abstract
Deductive reasoning is the process of deriving conclusions strictly from the given premises, without relying on external knowledge. We define honesty in this setting as a model’s ability to respond only when the conclusion is logically entailed by the premises, and to abstain otherwise. However, current language models often fail to reason honestly, producing unwarranted answers when the input is insufficient. To study this challenge, we formulate honest deductive reasoning as multi-step tasks where models must either derive the correct conclusion or abstain. We curate two datasets from graph structures, one for linear algebra and one for logical inference, and introduce unanswerable cases by randomly perturbing an edge in half of the instances. We find that prompting and existing training methods, including GRPO with or without supervised fine-tuning initialization, struggle on these tasks. In particular, GRPO optimize only for final task outcomes, leaving models vulnerable to collapse when negative rewards dominate early training. To address this, we propose Anchor, a reinforcement learning method that injects ground truth trajectories into rollouts, preventing early training collapse. Our results demonstrate that this method stabilizes learning and significantly improves the overall reasoning performance, underscoring the importance of training dynamics for enabling honest deductive reasoning in language models.
@@ 1 Introduction
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, their increasing deployment in real-world applications introduces critical safety considerations (betley2025emergent; raza2025industrial; bengio2025superintelligent; cloud2025subliminal). For these models to be deployed reliably, it is not sufficient for them to be merely helpful and harmless. They must also be honest (askell2021general; greenblatt2024alignmentfakinglargelanguage; sheshadri2025languagemodelsfakealignment): they should both (1) be aware of their own knowledge boundaries and (2) recognize whether a question is answerable from the information provided, to avoid fabricating information (jiang2021can; yin2023large; mohri2024language; kalai2025languagemodelshallucinate). However, existing benchmarks overwhelmingly focus on the first dimension, particularly the acknowledgement of factual uncertainty and knowledge boundaries (joshi2017triviaqa; kwiatkowski-etal-2019-natural; li2023halueval; niu2023ragtruth; yang2024alignment; guan2024hallusionbench), leaving the second dimension underexplored.
Deductive reasoning is a paradigm where the answerability of a conclusion depends solely on whether it can be derived from the premises stated in the prompt (clark1969linguistic). It offers a clean setting by isolating reasoning ability from factual recall, and allows us to define honest deductive reasoning as the behavior of producing a conclusion only when a valid derivation exists, and abstaining otherwise.
@@
Several training approaches have been widely used to enable models to perform reasoning tasks. Supervised fine-tuning (SFT) (wu2025generalization) has proven highly effective at quickly aligning models to desired behaviors, but it tends to overfit to demonstrations and struggles to generalize beyond the dataset distribution. Reinforcement learning methods such as Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) compares multiple rollouts of the same query to assign relative advantages and optimizes only for final verifiable outcomes. However, when all rollouts in the batch are incorrect and receive identical rewards, their relative advantages collapse to zero, leading to several issues including vanishing gradients and reinforcing dishonest overconfidence (liu2025understanding; yu2025dapo; zheng2025group). To address these issues, studies have explored integrating SFT with RL. However, most either lack access to a verifiable ground-truth trajectory for guidance (yan2025learning; liu2025ghpo; huang2025blending; nath2025adaptive), or do not directly address the stability of policy gradient updates, leaving them still susceptible to zero-variance issues at certain training stages (wu2025thought; chen2025step; zhang2025policy). This motivates the need for new datasets and methods that specifically target honest deductive reasoning.
@@
To study honesty in deductive reasoning in a controlled setting, we construct two multi-step reasoning datasets in which each query is either answerable or unanswerable given the premises, providing a precise testbed for evaluating whether models can reason honestly and recognize when valid reasoning paths exist and when they do not. The first dataset, GraphLA, is grounded in linear algebra: queries correspond to solving systems of equations along reasoning paths, while unanswerable cases are created by perturbing the system so that no valid solution path exists. The second dataset, GraphLI, is based on logical inference: queries test whether a conclusion follows from composed chains of implications, with unanswerable instances generated by removing or altering key premises or conclusions. Based on these datasets, we investigate the following two central research questions:
@@
For evaluation (RQ1), we show that across three model scales, performance on our benchmarks sharply declines as reasoning depth and problem size increase. Models struggle not only to follow reasoning chains but also to refrain from producing unwarranted conclusions when no valid derivation exists, revealing a lack of honest reasoning. For training (RQ2), we find that standard SFT and GRPO fail to overcome these challenges. Curriculum learning, when easy datasets are carefully and properly constructed, achieves strong performance but remains fragile and highly sensitive to difficulty calibration. In contrast, Anchor consistently stabilizes reinforcement learning, achieving robust performance on both answerable and unanswerable queries. When paired with curriculum learning, Anchor provides further gains, underscoring its strength in guiding models toward stable and honest deductive reasoning.
@@
    We formalize honesty in deductive reasoning as the ability to abstain on unanswerable queries and answer correctly on answerable queries, and introduce two datasets, GraphLA and GraphLI, that balance answerable and unanswerable cases.
@@
2 Related Work
Honesty Alignment
askell2021general define alignment via the “HHH” principles: helpful, honest, harmless. Honesty is an overloaded term, but it involves two key dimensions that help prevent models from fabricating information: (i) recognizing their own limitations, such as lacking the necessary knowledge or confidence; and (ii) recognizing whether a question is answerable from the available clues. Existing benchmarks that study honesty often conflate these two dimensions (joshi2017triviaqa; kwiatkowski-etal-2019-natural; li2023halueval; niu2023ragtruth; guan2024hallusionbench), making it difficult to isolate honesty in sense (ii) (yin2023large; ouyang2025treecut). kirichenko2025abstentionbench directly target this aspect, and our datasets complement their work by focusing on deductive reasoning without external knowledge, cleanly separating (ii) from (i). kalai2025language argue that hallucinations arise systemically and advocate evaluation standards that treat “I don’t know” as a strength. Our work aligns with this by emphasizing honest reasoning on tasks requiring recognition of unanswerable queries. [^1] See Appendix A for clarifications of a list of concepts.
Stabilizing Reinforcement Learning
SFT (behavior cloning) can be viewed as a special case of RL (wu2025generalization), where the policy imitates expert trajectories without exploration. Modern policy gradient methods often pair stochastic policies with advantage estimation to improve learning stability (williams1992simple; schulman2015high; schulman2017proximal). GRPO (shao2024deepseekmath) removes the value function and computes advantages in a group-relative manner. Despite its simplicity, GRPO exhibits biased optimization toward longer responses, struggles with overly easy or hard instances (liu2025understanding), suffers entropy collapse under poor exploration (yu2025dapo), and produces noisy gradients due to token-level importance ratios (zheng2025group).
Anchor specifically addresses GRPO’s failure on overly difficult queries by adding an SFT-like objective that anchors learning when exploration fails. Unlike prior approaches, it combines demonstrations with policy rollouts while avoiding SFT overfitting. Deep Q-learning from Demonstrations (hester2018deep) incorporates demonstrations via replay buffers, whereas Anchor injects them deterministically into each rollout group, which is advantageous for reasoning tasks where ground truth is available but exploration collapses. PSFT (zhu2025proximal) constrains policy updates during imitation, while Anchor aligns supervised and reinforcement signals more dynamically within each update. Concurrently, chen2025beyond unify SFT and RL via bilevel optimization, though progress may stall without reward signals. Other recent works combine SFT with RL (yan2025learning; liu2025ghpo; huang2025blending; nath2025adaptive; wu2025thought; chen2025step; zhang2025policy), but generally lack verifiable ground-truth trajectories or do not address instability in policy gradient updates.
@@ 4 RQ1 (Motivation): How Do Untrained Models Perform on Reasoning Tasks of Varying Deductive Difficulty?
We evaluate three open-sourced models: Qwen-2.5-3B-Instruct, Qwen-3-0.6B, and Qwen-3-1.7B. Experiment setup details are provided in § D.1.
@@ 6 Experiments
We experiment with the following baselines and approaches: Chain-of-Thought (CoT) prompting (wei2022chain), supervised fine-tuning (SFT), GRPO (shao2024deepseekmath), SFT+GRPO (where SFT is used as a cold start followed by GRPO), and Easy-to-Hard curriculum learning (where GRPO is first trained on an easier dataset and then on the target dataset). We report overall accuracy, accuracy on the unanswerable subset, and accuracy on the answerable subset.
@@
For GraphLI, we fix the reasoning depth to $k=15$ and the number of irrelevant edges to $|E_{\text{irr}}|=5$. For each configuration, 3 examples are sampled for instantiation. The dataset is split into 5316 training (2702 answerable, 2614 unanswerable), 300 validation (143 answerable, 157 unanswerable), and 300 test (155 answerable, 145 unanswerable) examples. For the easy dataset in curriculum learning, we reduce the reasoning depth to $k\in\{2,3,4,5\}$. Additional experimental details are provided in Appendix G.
(a) Qwen-2.5-3B-Instruct
(b) Qwen-3-1.7B
@@ 8 Conclusion
We investigated the challenge of aligning reasoning language models with honesty, focusing on tasks that require both solving answerable queries and abstaining on unanswerable ones. Our analysis showed that existing approaches such as SFT and GRPO either fail to provide reliable learning signals or collapse when faced with uniformly negative rewards. We proposed Anchor, a ground-truth–injected reinforcement learning method that stabilizes training by ensuring positive reference signals during rollouts. Across both the GraphLA and GraphLI datasets and multiple models, Anchor consistently outperformed baselines and proved robust where curriculum learning was fragile. Moreover, Anchor integrates seamlessly with Easy-to-Hard training, yielding further gains. These results show a step toward steadier reinforcement learning that enables honest deductive reasoning in language models.
@@
We have taken several steps to ensure the reproducibility of our work. A detailed description of dataset construction and statistics is provided in §§ 4 and 6, along with complete methodology details in § 5. Parameters and hyperparameters used for training and evaluation are documented in §§ 6, D.1 and G. Theoretical guarantees for our proposed method are supported by a formal proof of Proposition 1 in Appendix F. To further facilitate reproducibility, we will release both the datasets and source code upon acceptance.
Appendix A Concept Clarifications
As recent papers often overload terms such as honesty with different usages, leading to potential confusion, we provide our interpretations in Table 2. Our goal is to ensure clarity about how these terms are used in this paper and to avoid misunderstandings.
Concepts Definition
Honesty Honesty means that models should not fabricate information (askell2021general). This has two dimensions when a language model generates an answer. First, if the question is valid but challenging, the model should acknowledge its own knowledge boundaries or limitations. Second, if the question is invalid, the model should point this out rather than making up an answer. Most existing papers focus on the first dimension (yang2024alignment). We advocate for greater attention to the second dimension.
Abstention Abstention means choosing not to give an answer or make a decision, similar in meaning to refusal (madhusudhan2025llms; wen2025know). Some papers use the term in a narrower way. For example, kirichenko2025abstentionbench describe abstention as a response in which a model explicitly acknowledges an issue in the user’s question. However, this interpretation narrows the original meaning. Models may also abstain for safety reasons when a query is harmful.
Factuality Factuality refers to how well a model’s output aligns with ground truth world knowledge and is often used interchangeably with truthfulness. This fits within the first dimension of honesty but emphasizes the complementary case that when the model does know the answer it should state it correctly.
Deductive Reasoning Deductive reasoning is a process in which the conclusion is guaranteed to be true if all premises are true and the inference rules are followed. It requires that all information needed for the conclusion is provided explicitly in the input rather than relying on implicit external knowledge.
Table 2: Clarifications of a list of concepts.
@@ E.1 GRPO
All tokens within a rollout $y_{i}$ share the same normalized advantage $\hat{A}_{i}$. The corresponding policy gradient is
@@ E.2 SFT
The SFT objective gradient is simply the logarithmic likelihood gradient on the supervised dataset, with no advantage weighting:
@@ Appendix G Experiment Setup
Appendix H Additional Discussion on Process Rewards
We also extensively explored a wide range of process reward designs and found none to be effective. We experimented with several settings:
  1. 1.
    extracting intermediate reasoning steps from trajectories using sentence boundaries, newline characters, or explicit markup such as <step></step> tags;
  2. 2.
    assigning outcome rewards only to tokens within <answer></answer> while giving separate process rewards to tokens within <think></think>, or combining outcome and process rewards for pre-answer tokens using a weighted formulation; and
  3. 3.
    generating process rewards using LLM-as-a-judge signals, rule-based matching of intermediate variable values under explicit instructions, and entropy-based heuristics.
Across all three model sizes and both datasets, these attempts failed to provide meaningful learning signals. We found that designing appropriate process rewards for each reasoning step was extremely challenging, and even when a plausible reward signal existed, it was highly task-specific (e.g., differing substantially between linear algebra and logical inference). This task specificity runs counter to our goal of developing a generally applicable training method.
Appendix I Clarifications on LLM Usage